OpenStudy (anonymous):

will give medal & fan!!(:

OpenStudy (anonymous):

@ryan123345 @satellite73

OpenStudy (anonymous):

@Chelsea04

OpenStudy (anonymous):

for the first one, the center isn't even in the fourth quadrant, so ignore that one

OpenStudy (anonymous):

do you know how to check where each centre is?

OpenStudy (anonymous):

for the second one, the center is \((5,-7)\) and the radius is \(\sqrt{16}=4\) so that one stays completely inside the 4th quadrant, because 4 is less than 5 and 7

OpenStudy (anonymous):

can you check the other two?

OpenStudy (anonymous):

do you know how to check?

OpenStudy (anonymous):

No how do you? :) @Chelsea04

OpenStudy (anonymous):

ok, the formula for circles is: (x-h)^2+(x-k)^2=r^2

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

do you know what the center of the third one is? use \[(x-h)^2+(y-k)^2=r^2\] for a circle with center \((h,k)\) and radius \(r\)

OpenStudy (anonymous):

(h,k) are the coordinates for your centre. r is the radius of your centre

OpenStudy (ryan123345):

ill help him lol...

OpenStudy (anonymous):

so for the 3rd one it's -4 for h And 2 for k?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

sorry, the opposite signs, because when you take it out of the brackets, it changes sign

OpenStudy (anonymous):

yeah! my bad @Chelsea04

OpenStudy (anonymous):

so the coordinates for your centre is (4,-2)

OpenStudy (anonymous):

yeah is that a 4th quad? yeah right? :)

OpenStudy (anonymous):

yes, but the question asks for the circle to be completely in the 4th quad

OpenStudy (anonymous):

oh and it's like only half! lol so I got B and D (:

OpenStudy (anonymous):

so, now is where the radius comes into the question. so remember the equation used r^2, so you have to remember to square root that number :)

OpenStudy (anonymous):

but yes, those would be your answers

OpenStudy (anonymous):

do you understand?